# Status Report on the Chicago-Carnegie Hubble Program (CCHP): Three Independent Astrophysical Determinations of the Hubble Constant Using the James Webb Space Telescope

### v1

@@ 6.2 JAGB Stars
Although relatively new in the context of the extragalactic distance scale, J-region AGB (JAGB) stars were, in fact, first identified in the LMC nearly a quarter of a century ago, as a distinct class of objects (Nikolaev & Weinberg, 2000; Weinberg & Nikolaev, 2001), when they were used to determine the back-to-front geometry of the LMC. The JAGB method is now one of the most promising methods for measuring the distances to nearby galaxies. (1) These thermally-pulsating AGB carbon stars have a nearly constant luminosity in the near-infrared J band (at 1.2 m), and (2) they have a low intrinsic dispersion of only 0.2 mag Nikolaev & Weinberg (2000); Madore & Freedman (2020). Moreover, (3) JAGB stars can be easily and unambiguously identified on the basis of their near-infrared colors alone (without the need for spectroscopy or narrow-band photometry), thereby being readily distinguished from both the (bluer) O-rich AGB stars, as well as (even redder) “extreme” carbon stars.
Extending the JAGB method, Freedman & Madore (2020) used these stars to measure the distances to additional nearby galaxies within and beyond the Local Group, out to 27 Mpc. In a sample of 14 galaxies, they found excellent agreement of the JAGB distances with published TRGB distances to those same galaxies, where the combined scatter (including potential effects of metallicity and star formation history) amounts to only 4%. Simultaneously, Ripoche et al. (2020) investigated the Magellanic Clouds and the Milky Way. A number of additional extensive tests of this method have recently been carried out by Lee and collaborators Lee et al. (2021, 2022); Lee (2023); Lee et al. (2024a, b) as well as Parada et al. (2021) and Zgirski et al. (2021) in several nearby galaxies, confirming the excellent agreement with distances measured with the TRGB and Cepheid distance scales, and again indicating that any potentially confounding effects of metallicity variations and star formation history must be contained within that small scatter, and therefore must be relatively small, individually and collectively.
@@
The advantages of the JAGB method are many. (1) JAGB stars are the dominant population of the reddest stars found in galaxies. They are easily identified on the basis of their near-infrared colors. (2) The color-selected -band luminosity function is centrally peaked with a low dispersion. (3) JAGB stars are, in the mean, about one magnitude brighter than the brightest TRGB stars. (4) JAGB stars are found in all galaxies that have an intermediate-age population, and therefore the method can be applied to a wide range of galaxy types. (5) In the infrared, the total reddening along the line of sight (including Milky Way foreground extinction, host galaxy extinction, and any dust reddening generated by the carbon stars themselves) drops by about a factor of four in going from the optical () to the near infrared (). And, (6) unlike the use of period-luminosity relations for variable stars (e.g., Cepheids or Miras), which need to be monitored over multiple cycles, JAGB stars require only a single-epoch observation.
@@
6.3 Cepheids
Cepheid variables have played a key role in the determination of extragalactic distances: over the last century, they became the de facto gold standard. For recent reviews of the Cepheid distance scale see Madore & Freedman (1991); Bono et al. (2010); Freedman & Madore (2010); Turner (2012); Freedman & Madore (2023a). Cepheid variability results from pulsation of their atmospheres, which is driven by a valve-like mechanism as helium cycles from a singly to a doubly ionized state, and the opacity increases as the star is compressed. The well-defined relationship between period, luminosity and color provides a standardizable candle of high precision. Including a metallicity term, the Leavitt law can be linearly parameterized as
where the Cepheid magnitude at a given wavelength is a function of the logarithm of the period (P) with coefficient , a color term with coefficient , and a term with coefficient , that allows for a metallicity dependence (where [O/H] represents the logarithmic oxygen-to-hydrogen ratio for HII regions in the vicinity of the Cepheids, relative to the solar value); and is the overall zero point.
A well-known characteristic of the PL relation is its decreasing scatter/width with increasing wavelength (e.g., Madore & Freedman, 1991). The simultaneous decrease of both reddening and metallicity effects, again with increasing wavelength, motivates the application of the Leavitt Law at longer (near- and mid-infrared) wavelengths (McGonegal et al., 1982; Madore & Freedman, 1991; Freedman et al., 1991; Macri et al., 2001; Freedman et al., 2008). (An exception is the 4.5m band in which the Cepheid flux is affected by the presence of a CO bandhead (Scowcroft et al., 2011)). Hence, the focus of recent applications to the Cepheid distance scale has been to discover these variables at optical wavelengths, and then following them up with available redder passbands (e.g., Freedman et al., 2001; Macri et al., 2001; Freedman et al., 2012; Riess et al., 2009, 2022).
If accurate colors are available, a reddening-free magnitude or Wesenheit function, , can be constructed, as first demonstrated by Madore (1976, 1982). is widely used in applications of the Cepheid distance scale (e.g., Freedman et al., 2001; Riess et al., 2016). For example, at optical wavelengths
where RV = AV /E(B-V) is the ratio of total-to-selective absorption. simultaneously corrects for total line-of-sight absorption, including both host-galaxy (internal) and Galactic (foreground) contributions.
The advantages of the Cepheid distance scale are many. (1) Cepheids are supergiant stars placing them among the brightest stellar indicators. (2) Cepheids are relatively young, high-mass stars, found in abundance in star-forming galaxies. (3) The Cepheid Leavitt relation, at the reddest wavelengths, has a small scatter (e.g., in the I or J band), where the dispersions amount to only 0.1 to 0.2 mag. (4) Cepheids are observationally straightforward to discover and uniquely identify. At optical wavelengths, they exhibit amplitudes up to 1 mag, and they display characteristic and identifiable (“sawtooth”) light curve shapes. And, (5) Cepheids can be re-observed over time (unlike supernovae), thereby allowing for the updating of their periods and phases, and adding additional wavelength coverage or achieving greater image resolution as new instrumentation becomes available, as, for example, in the case of HST and JWST.
There are also disadvantages of the Cepheid distance scale. (1) Cepheid variables are generally found in the relatively high-surface-brightness areas in the star-forming disks of late-type galaxies. In these regions crowding/blending effects increase the random and systematic noise seen in the individual Cepheid measurements and complicate sky subtraction. (2) Being relatively young stellar objects, Cepheids are located close to the regions of dust and gas out of which they formed. Hence, precise and accurate colors are required in order to correct the stars individually for the effects of reddening and extinction. (3) The effects of metallicity on the Cepheid Leavitt Law as a function of wavelength are still being actively debated in the literature (e.g. Ripepi et al., 2020; Madore & Freedman, 2024a). (4) Complexity. The many steps required to standardize Cepheid luminosities (measurement of light curves, noise threshold for inclusion in sample, completeness bias, adopted period range, artificial star corrections to correct for crowding/blending effects, extinction corrections, metallicity corrections) all involve choices. Cepheids are not simple standard candles.
@@ 7 New Distances from JWST Observations
As noted in §3, our observing program was configured to optimize simultaneous observations for all three of our methods555The JWST data for our program are available from the Barbara A. Mikulski Archive for Space Telescopes (MAST) Portal, Proposal ID 1995.. This goal led us to choose as our primary filter. First, JAGB stars have nearly constant luminosities in the () band. Second, observations provide an additional waveband beyond the HST and filters, to improve the extinction correction for Cepheids. Third, extinction effects are small at 1.15 m, which is an advantage for the TRGB method relative to the band (where extinction effects are about a factor of two larger). We note that the disadvantage of the -band for the TRGB method is that there is a steeper color dependence than for the band where the TRGB is relatively flat (e.g., Lee et al., 1993). Nonetheless, the TRGB is still standardizable in the near-infrared (e.g., Dalcanton et al., 2012; Madore et al., 2018; Hoyt et al., 2018; Durbin et al., 2020; Hoyt et al., 2024).
As also noted in §3, our original choice for our second filter was motivated to allow us to provide constraints on the metallicity sensitivity of the Leavitt law. At 4.5 m there is a metallicity dependence of the Leavitt law as a result of a CO-band head (Scowcroft et al., 2011). The filter can also provide a broad color baseline for the JAGB method. However, the relatively low resolution of the data prompted us to switch to the filter, which provides both higher resolution and higher sensitivity.
@@ 7.1 TRGB Measurements
An important aspect of an accurate measurement of an old TRGB population is the spatial selection – finding an optimal balance between avoiding regions of the disk where younger populations reside, and not being so far out in the halo so that the number density of stars is too small. As described in H24, we used the blue, unresolved light to trace regions dominated by young disk populations, masking regions bright in blue flux.
@@
In Figure 4, we show examples of versus ( - ) color-magnitude diagrams for three of our program galaxies, NGC 4258, NGC 4424 and NGC 4039 from H24. The position of the TRGB is shown, and is (both qualitatively and quantitatively) easily identified. As can be seen, the TRGB feature in the infrared exhibits a slope with photometric color. This dependence is due to a combination of metallicity and, to a lesser extent, to age effects. In the band, the metallicity dependence is about five times larger than that due to age (e.g., McQuinn et al., 2019). To measure the TRGB in each SN Ia host, the gradient in the magnitude direction was computed from a smoothed Hess diagram of the selected RGB stars, and the centroid of the dominant contour defined the TRGB in each field.
Figure 5 illustrates a smoothed Hess diagram of RGB stars in the galaxy NGC 4424. For each galaxy, the magnitude-direction gradient contained a maximal set of contours that indicated the location of the TRGB. A threshold between the 90th and 99th percentile of the 2-D gradient was then used to define the TRGB contour. An ellipse was then fit to this TRGB contour and the centroid of that ellipse was adopted as the best-fit magnitude and color of the TRGB. To compute the measurement uncertainty, the length of the axis perpendicular to the sloped TRGB was scaled by the contrast of the TRGB ellipse, which was computed as the ratio of the number of stars contained on the faint end of the semi-major axis chord to those above it. The mean TRGB color and magnitude estimated for each SN Host and for three fields in NGC 4258 were then combined with an estimate of the TRGB color slope (see H24 for details) to determine distance moduli. With a mean TRGB color and magnitude estimated for each SN Ia host, as well as for each of three fields in NGC 4258, distance moduli were computed relative to NGC 4258.
@@ 7.2 JAGB Measurements
In brief, as described in detail in L24, the JAGB stars in our target galaxies were selected based on their position in (initially blinded) near-infrared () or () JWST NIRCam color-magnitude diagrams. The magnitudes were binned and the luminosity functions were smoothed using a Gaussian-windowed, Locally Weighted Scatterplot Smoothing (GLOESS) algorithm (e.g., Persson et al., 2004). The JAGB magnitudes for each galaxy were determined from the mode of the smoothed luminosity function.
@@
The JAGB method is best applied in the outer disks and halos of galaxies where there are sufficient numbers of carbon stars to provide a statistically meaningful measurement, but not too far into the disk where systematic effects from crowding, blending, and reddening become an issue, and the luminosity function becomes asymmetric. As demonstrated in L24, the JAGB luminosity exhibits a well-defined mode and a distinct Gaussian form in the lower-reddening, less crowded regions of the galaxies. (The mode is also robust against asymmetries in the underlying distribution, as well as robust against outliers in the tails and/or windowing or clipping.) The value of the mode is generally found to be brightest in the inner, high-surface brightness regions of a galaxy where the crowding and reddening effects are greatest. For a clean measurement of the JAGB luminosity function, these high surface-brightness regions need to be avoided.
@@
The radial cuts for the JAGB luminosity functions were set by seeking convergence in the radial distribution: that is, the radial distance, r, within the disk where the derivative of the magnitude of the JAGB (d/dr) stabilized and leveled off to zero (L24). In three cases (NGC 3972, NGC 4424 and NGC 4038), no clean convergence was found, and these galaxies were not included in further analysis.
7.3 Tests for Systematics in Distances Measured Using the JAGB Method
@@
To test how our choice of JAGB statistic affected the final measured distances, we explored how the smoothing parameter affected the final measured mode, since the mode measured from an increasingly smoothed luminosity function will eventually converge to the mean. We varied the smoothing scale of the JAGB luminosity function using smoothing scales of (0.15, 0.20, 0.25, 0.30, 0.35, 0.40) mag, and then re-measured the mode with the given smoothing scale. We defined the statistical error due to the choice of smoothing parameter as the maximum difference between the fiducial mode (measured with a smoothing parameter of mag for all galaxies) and all of the measured modes. The total smoothing parameter error was then defined to be the smoothing parameter error from the SN Ia host galaxy and the zeropoint from NGC 4258 added in quadrature. This error was larger than 0.05 mag for all galaxies. In Lee+ (2024, submitted), we demonstrated that using the mean/median instead of the mode as the chosen JAGB statistic resulted in distance moduli that were measured to be 0.03 mag brighter on average. Thus, this systematic offset was fully encapsulated within the minimum adopted smoothing parameter uncertainty of 0.05 mag. Differences between the mean and the modal values of the JAGB populations were also explored in the study of Madore et al. (2022); see their Figure 1.
@@
7.4 Cepheid Measurements
At present, measurement of the Cepheid distances cannot be obtained from our JWST data alone: to correct for extinction necessitates measurements at multiple wavelengths. For this purpose we used existing HST/ archival data at optical ( and ) wavelengths obtained by the SHoES team. We independently reduced all of the archival data for our target galaxies.
Initial Cepheid candidates were obtained directly from Riess et al. (2016, 2022). Periods and mean magnitudes for these candidates were obtained using template light curves from O24. We eliminated candidates with poor light curve fits or extremely low amplitudes (i.e., 0.2 mag in the or band: the filter used was that upon which the published periods were based), as well as candidates with periods in excess of 100 days. Examples of PL relations at , , and NIRCam are shown in Figure 6 for the galaxies NGC 5643 and NGC 4639. In order to correct the magnitudes for crowding bias, we carefully injected 1,000 artificial stars, one at a time, within a 3 arcsec radius of each Cepheid. Following the results of the artificial star tests, we eliminated Cepheids that were close to the edge of the NIRCam frames, as they had lower quality photometry due to non-overlapping dithers. We fit period-luminosity (PL) relations using fixed slopes from Monson et al. (2012) in the NIRCam , , , and bands. Finally, we excluded obvious outliers based on a 4-sigma clip.
For our primary analysis, we calculated PL relation intercepts (and their uncertainties) through resampling, including redrawing from the distribution of the artificial star corrections for each Cepheid. The PL relation intercepts and uncertainties were taken to be the mean and standard deviation of the simulated distribution. We fit a Cardelli et al. (1989) reddening law with Rv = 3.1 to the set of resampled PL relations, extrapolating to infinite wavelength, and adopted the mean and standard deviation of the distribution of reddening law fits as the absolute distance modulus and statistical uncertainty. In addition, we carried out two additional cross-checks on the distances. We also directly measured PL relations and reddening law fits without weighting or resampling, adopting the mean artificial star corrections for each Cepheid and for each bandpass. In addition, we constructed Wesenheit PL relations (Madore, 1976) using the magnitudes and () colors to measure unweighted reddening-corrected distances. The median offset of the absolute distance moduli compared to the primary analysis for the 10 SN Ia host galaxies was 0.01 mag for both comparisons.
Following on from our blind analysis, we also undertook a broader analysis of the fitting of the Cepheid PL relations. In this case, we employed a Sobel edge-detector to identify the upper and lower 2- edges of the instability strip that define the boundaries of the PL relations. This was done by marginalizing the PL relations using previously determined fixed slopes. The intrinsic widths of the PL relations as a function of wavelength were obtained from well-observed Cepheids in the LMC. This method was found to be quite insensitive to choice of period cuts and incomplete sampling within the instability strip. It also provided a means of identifying outliers. It yielded significantly larger distance moduli, primarily a function of the fact that within our calibrating galaxy, NGC 4258, there are few faint stars with short periods, as well as few bright Cepheids with longer periods, thereby leading to a significant flattening of the PL relations due to the underpopulated extremes of the PL relations.
One can, of course, make a series of choices in how to define the Cepheid sample selection, whether to impose period cuts, how to measure the extinction, and what, if any, metallicity correction is required. As our independent analyses have revealed, one can make reasonable (defensible) choices, but obtain significantly different results. Moreover, it is non-trivial to quantitatively assess the impact of these choices on the error budget. Unfortunately, these subtle data-reduction choices can determine whether or not there is an tension. Our view is that these discrepant values provide valuable insight into the current uncertainties. We leave to future studies (with improved colors, and smaller scatter in the PL relations) a resolution of these discrepancies. Because running our PL fits with various period cuts leads to significantly different distance moduli, at this juncture, we conservatively adopt an additional uncertainty of 0.07 mag for our Cepheid distances.
8 Comparison of the Cepheid, TRGB and JAGB Distances

### v2

@@ 6.2 JAGB Stars
Although relatively new in the context of the extragalactic distance scale, J-region AGB (JAGB) stars were, in fact, first identified in the LMC nearly a quarter of a century ago, as a distinct class of objects (Nikolaev & Weinberg, 2000; Weinberg & Nikolaev, 2001), when they were used to determine the back-to-front geometry of the LMC. The JAGB method is now one of the most promising methods for measuring the distances to nearby galaxies. (1) These thermally-pulsating AGB carbon stars have a nearly constant luminosity in the near-infrared J band (at 1.2 m), and (2) they have a low intrinsic dispersion of only 0.2 mag (Nikolaev & Weinberg (2000); Madore & Freedman (2020)). Moreover, (3) JAGB stars can be easily and unambiguously identified solely on the basis of their near-infrared colors alone (without the need for spectroscopy or narrow-band photometry), thereby being readily distinguished from both the (bluer) O-rich AGB stars, as well as (even redder) “extreme” carbon stars.
Extending the JAGB method, Freedman & Madore (2020) used these stars to measure the distances to additional nearby galaxies within and beyond the Local Group, out to 27 Mpc. In a sample of 14 galaxies, they found excellent agreement of the JAGB distances with published TRGB distances to those same galaxies, where the combined scatter (including potential effects of metallicity and star formation history) amounted to only 4%. Simultaneously, Ripoche et al. (2020) investigated the Magellanic Clouds and the Milky Way. A number of additional extensive tests of this method have recently been carried out by Lee and collaborators Lee et al. (2021, 2022); Lee (2023); Lee et al. (2024a, c) as well as Parada et al. (2021), Zgirski et al. (2021), and Magnus et al. (2024) in several nearby galaxies, confirming the excellent agreement with distances measured with the TRGB and Cepheid distance scales, and again indicating that any potentially confounding effects of metallicity variations and star formation history must be contained within that small scatter, and therefore must be relatively small.
@@
The advantages of the JAGB method are many. (1) JAGB stars are the dominant population of the reddest stars found in galaxies. They are easily identified on the basis of their near-infrared colors. (2) The color-selected -band luminosity function is centrally peaked with a low dispersion. (3) JAGB stars are, in the mean, about one magnitude brighter than the brightest TRGB stars. (4) JAGB stars are found in all galaxies that have an intermediate-age population, and therefore the method can be applied to a wide range of galaxy types. (5) In the infrared, the total reddening along the line of sight (including Milky Way foreground extinction, host galaxy extinction, and any dust reddening generated by the carbon stars themselves) drops by about a factor of four in going from the optical () to the near infrared (). (6) Another major advantage is that JAGB stars can be found in the outer disks of galaxies where crowding and reddening effects are significantly reduced compared to the inner disks where Cepheids reside. And, (7) unlike the use of period-luminosity relations for variable stars (e.g., Cepheids or Miras), which need to be monitored over multiple cycles, JAGB stars require only a single-epoch observation.
@@ 7 New Distances from JWST Observations
As noted in §3, our observing program was configured to optimize simultaneous observations for all three of our methods555The JWST data for our program are available from the Barbara A. Mikulski Archive for Space Telescopes (MAST) Portal, Proposal ID 1995.. This goal led us to choose as our primary filter. First, JAGB stars have nearly constant luminosities in the () band. Second, observations provide an additional waveband beyond the HST and filters, to improve the extinction correction for Cepheids. Third, extinction effects are small at 1.15 m, which is an advantage for the TRGB method relative to the band (where extinction effects are about a factor of two larger). We note that the disadvantage of the -band for the TRGB method is that there is a steeper color dependence than for the band where the TRGB is relatively flat (e.g., Lee et al., 1993). Nonetheless, the TRGB is still standardizable in the near-infrared (e.g., Dalcanton et al., 2012; Madore et al., 2018; Hoyt et al., 2018; Durbin et al., 2020; Newman et al., 2024; Hoyt et al., 2024). Solving for a slope makes the method more analogous to that of measuring a period-luminosity relation for Cepheids.
As also noted in §3, our original choice for our second filter was motivated to allow us to provide constraints on the metallicity sensitivity of the Leavitt law. At 4.5 m there is a metallicity dependence of the Leavitt law as a result of a CO-band head (Scowcroft et al., 2011). The filter would also provide a broad color baseline for the JAGB method. However, the relatively inferior resolution of the imaging data prompted us to switch to the filter, which provides both higher resolution and higher sensitivity.
@@ 7.1 TRGB Measurements
An important aspect of accurately measuring TRGB distances is in selecting for an old population. Doing so requires an optimal balance between avoiding regions of the disk where younger populations reside, and not being so far out in the halo so that the number density of stars is too small (for a fixed-sized detector footprint on the sky). As described in H25, we used the unresolved blue light to trace regions dominated by young disk populations, masking those regions bright in blue flux and with noisy, unclear TRGBs.
@@
In Figure 5, we show examples of versus ( - ) color-magnitude diagrams for three of our program galaxies, NGC 4258, NGC 4424 and NGC 4039 from H25. The position of the TRGB is shown, and is (both qualitatively and quantitatively) easily identified. As can be seen, the TRGB feature in the infrared exhibits the expected upward trend with increasing photometric color. This dependence is due to a combination of metallicity and, to a much lesser extent, age effects (e.g., McQuinn et al., 2019). To measure the TRGB in each SN Ia host, the gradient in the magnitude direction was computed from a smoothed Hess diagram of the selected RGB stars, and the centroid of the dominant contour defined the TRGB in each field. The different fields show some variation in slope, as seen here, but overall there is a consistent single, universal slope plus dispersion, all of which is propagated into the TRGB distance error budget (see Hoyt et al., 2025, for further details).
7.2 TRGB Measurements
Archival JWST data from the SHoES Cycle 1 program (GO-1685 , PI: A. Riess) offer an important constraint on our TRGB calibration, given that they were taken with different filters that are complementary to our own observations. The galaxies observed include NGC 1448, NGC 1559, NGC 4258, NGC 5584 and NGC 5643. The data have been reduced in a similar fashion to those described in our previous optical observations (e.g., Freedman et al., 2019; Hoyt et al., 2019; Jang et al., 2021). Two members of our CCHP group reduced the data independently in blind fashion, finding agreement in the average at the 0.015 mag (0.7%) level.
The data for NGC 4258 were used to measure the zero point for the TRGB. In the left panel of Figure 6, we show an versus () color-magnitude diagram for a halo region of NGC 4258. The position of the TRGB is shown by the solid black horizontal line at F090W = 25.08 mag. The right panel shows the marginalized and smoothed luminosity function in black, and the Sobel edge-detection filter output is shown in blue. Detection of the tip in the left panel, and its quantitative measurement in the right panel, is unambiguous. The luminosity function was smoothed by 0.1 mag before its digital first derivative was output by the Sobel filter.
The JWST -band () data were derived from the sum of four 258-second NIRCam exposures; the four equivalent exposures were taken in parallel, directed to the beam-split long-wavelength channel of NIRCam. The Sobel response function has a measured 2- width of 0.20 mag. Within the 4 full width of 0.4 mag there are 624 red giant branch stars contributing to the measurement. We therefore adopt a value of 0.4/0.02 mag for the statistical uncertainty.
We adopt a foreground -band galactic line-of-sight extinction of 0.02 mag for the halo of NGC 4258, as derived by Anand et al. (2024a). This gives a reddening-corrected magnitude of 25.06 mag for the TRGB in NGC 4258. Adopting a geometric distance modulus of 29.397 0.032 mag (Reid et al., 2019) yields our zero-point calibration of MF090W = -4.336 0.02 (stat) 0.032 [sys] mag. Note this value is on the Vega-Sirius system, the default magnitude system employed by the JWST calibration program (Rieke et al., 2023; Gordon et al., 2022). The measured distance moduli are = 31.321 0.049, = 31.491 0.051, = 31.851 0.053, and = 30.599 0.057. R24b measured TRGB distances of = 31.38 0.07, = 31.50 0.05, = 31.81 0.09, and = 30.56 0.06. These distance moduli agree very well, with a weighted mean difference amounting to only -0.003 0.038 mag ( CCHP-SHoES). Two additional galaxies from this program are not yet publicly available, NGC 2525 and NGC 3447.
7.2.1 Crosschecks of the TRGB F090W Zero Point
Our zero-point value of MF090W = -4.336 0.02 (stat) 0.032 [sys] mag compares favorably with the zero point of MF090W = -4.32 0.025 (stat) mag published by Newman et al. (2024). They averaged over six nearby galaxies observed with JWST, using the same filter combinations as discussed here, but zeroed to previously published TRGB distances using an -band calibration of MI = -4.05 mag (Freedman, 2021). The SHoES team (Anand et al., 2024b) has also published an averaged zero point on the Vega-Vega system for the TRGB method, giving MF090W = -4.362 0.033 (stat) 0.045 [sys] mag. Accounting for the approximately 0.035 mag offset between the Vega-Vega and Vega-Sirius systems for , their value is -4.327 mag.
The full range of all three values spans 0.014 mag, suggesting excellent convergence toward an zero point. Furthermore, because the Newman et al. (2024) value was based on assuming an ingoing zero point equal to -4.05 mag, the remarkable consistency between all three measurements means the Newman et al., CCHP, and SHoES all agree on the TRGB’s zero point being equal to -4.05 mag. These comparisons provide a strong test of the accuracy of the zero point calibration (adopted from F21) that underlies the HST TRGB distances used in our primary analysis based on 24 calibrator SNe Ia.666 Note that if we adopt an -band TRGB zero-point of -3.95 mag, as given in many SHoES papers, the Newman et al. zero-point would be 0.1 mag fainter, becoming -4.22 (=-4.32 + 0.1). This value deviates significantly from the SHoES zero point (-4.36), at greater than a 2-sigma level.
@@ 7.3 JAGB Measurements
In brief, as described in detail in L25, the JAGB stars in our target galaxies were selected based on their position in (initially blinded) near-infrared () or () JWST NIRCam color-magnitude diagrams. The magnitudes were binned and the luminosity functions were smoothed using a Gaussian-windowed, Locally Weighted Scatterplot Smoothing (GLOESS) algorithm (e.g., Persson et al., 2004). The JAGB magnitudes for each galaxy were determined from the mode of the smoothed luminosity function.
@@
The JAGB method is best applied in the outer disks and halos of galaxies where there are sufficient numbers of carbon stars to provide a statistically meaningful measurement, but not too far into the disk where systematic effects from crowding, blending, and reddening become an issue, and the luminosity function becomes asymmetric. As demonstrated in L25, the JAGB luminosity exhibits a well-defined mode and a distinct Gaussian form in the lower-reddening, less-crowded regions of the galaxies. The mode is also robust against asymmetries in the underlying distribution, as well as robust against outliers in the tails and/or windowing or clipping. The value of the mode is generally found to be brightest in the inner, high-surface brightness regions of a galaxy where the crowding and reddening effects are greatest. For a clean measurement of the JAGB luminosity function, these high-surface-brightness regions need to be avoided.
@@
The radial cuts for the JAGB luminosity functions were set by seeking convergence in the radial distribution: that is, the radial distance, R, within the disk where the derivative of the magnitude of the JAGB (d/dR) stabilized and leveled off to zero (L25). In three cases (NGC 3972, NGC 4424 and NGC 4038), a clear convergence was not found, and these galaxies were not included in further analysis.
7.3.1 Tests for Systematics in Distances Measured Using the JAGB Method
@@
To test how our choice of JAGB statistic affected the final measured distances, we explored how the smoothing parameter affected the final measured mode, since the mode measured from an increasingly smoothed luminosity function will eventually converge toward the mean. We varied the smoothing scale of the JAGB luminosity function using smoothing scales of (0.15, 0.20, 0.25, 0.30, 0.35, 0.40) mag, and then re-measured the mode with the given smoothing scale. We defined the statistical error due to the choice of smoothing parameter as the maximum difference between the fiducial mode (measured with a smoothing parameter of mag for all galaxies) and all of the measured modes. The total smoothing parameter error was then defined to be the smoothing parameter error from the SN Ia host galaxy and the zeropoint from NGC 4258 added in quadrature. This error was larger than 0.05 mag for all galaxies. In Lee et al. (2024b), we demonstrated that using the mean/median instead of the mode as the chosen JAGB statistic resulted in distance moduli that were measured to be 0.03 mag brighter on average. Thus, this systematic offset was fully encapsulated within the minimum adopted smoothing parameter uncertainty of 0.05 mag. Differences between the mean and the modal values of the JAGB populations were also explored in the study of Madore et al. (2022a); see their Figure 1.
@@
8 Comparison of the TRGB and JAGB Distances
In Table 2, we list the supernovae, host galaxy name, individual distances () plus their uncertainties () and the weighted averages () and standard errors ) for both of the methods. Detailed descriptions of the measurement of these distances and their errors are provided in H25 (Tables 4 and 5) and L25 (Tables 3 and 5). For the averaging, individual weights are given by
where the weighted mean is
and the weighted standard deviation is
For each SN Ia-host distance average, we incorporated the respective zero-point measurement uncertainty for NGC 4258: 0.025 mag for the TRGB and 0.041 mag for the JAGB. These errors are correlated within each method’s distance scale, but uncorrelated between methods. Consequently, the combined uncertainty of 0.021 mag, calculated as . This uncertainty is correlated across all averaged TRGB and JAGB distances. Therefore, we removed it in quadrature from the values in Table 2, and added it back into the final error budget as a systematic uncertainty.